### v1

# Quality-aware Masked Diffusion Transformer for Enhanced Music Generation

Abstract
In recent years, diffusion-based text-to-music (TTM) generation has gained prominence, *Equal contribution.
Code on https://github.com/ivcylc/qa-mdt
offering a novel approach to synthesizing musical content from textual descriptions. Achieving high accuracy and diversity in this generation process requires extensive, high-quality data, which often constitutes only a fraction of available datasets. Within open-source datasets, the prevalence of issues like mislabeling, weak labeling, unlabeled data, and low-quality music waveform significantly hampers the development of music generation models. To overcome these challenges, we introduce a novel quality-aware masked diffusion transformer (QA-MDT) approach that enables generative models to discern the quality of input music waveform during training. Building on the unique properties of musical signals, we have adapted and implemented a MDT model for TTM task, while further unveiling its distinct capacity for quality control. Moreover, we address the issue of low-quality captions with a caption refinement data processing approach. Our demo page is shown in https://qa-mdt.github.io/.
1 Introduction
Figure 1: The distribution curves of CLAP similarity and pseudo-MOS for large-scale open-source music databases AudioSet (Defferrard et al., 2016) and FMA (Defferrard et al., 2016), where darker areas represent higher text-audio alignment or audio quality.
Text-to-Music (TTM) generation aims to transform textual descriptions of emotions, style, instruments, rhythm, and other aspects into corresponding music segments. This provides new expressive forms and innovative tools for multimedia creation. According to the scaling law principles (Peebles and Xie, 2023; Li et al., 2024), effective generative models require a large volume of training data. However, unlike previous image generation tasks  (Chen et al., 2024a; Rombach et al., 2021), acquiring high-quality music data often faces greater challenges. This is primarily due to copyright issues and the need for professional hardware equipment to capture high-quality music data. This makes building a high-performance TTM model particularly difficult.
In the TTM field, high-quality paired data of text and music signals is scarce, this prevalent issue of low-quality data, highlighted in Figure 1, manifests in two primary challenges. Firstly, there is a weak correlation between music signals and captions, characterized by missing, weak, or incorrect captions, leading to low text-audio similarity which can be indicated by CLAP score (Wu* et al., 2023). Secondly, most available music signals often suffer from distortion due to noise, low recording quality, or outdated recordings, resulting in diminished generated quality which can be measured by pseudo-MOS scores from quality assessment models (Ragano et al., 2023). These challenges significantly hinder the training of high-performance music generation models, leading to poor rhythm, noise, and inconsistencies with textual control conditions in the generated audio. Therefore, effectively training on large-scale datasets with label mismatches, missing labels, or low-quality waveform has become an urgent issue to address.
In this study, we introduce a novel quality-aware masked diffusion transformer (QA-MDT) to enhance music generation. This model effectively leverages extensive, open-source music databases, often containing data of varying quality, to produce high-quality and diverse music. During training, we inject music quality through quantified pseudo-MOS (p-MOS) scores into the denoising stage at multiple granularities to foster quality awareness, while coarse-level quality information is seamlessly integrated into the text encoder and fine-level details are embedded into the transformer based diffusion architecture. Masking strategy is also employed to enhance the spatial correlation of the music spectrum and further accelerates convergence. This innovative strategy guides the model in the generation phase to produce high-quality music by leveraging information corresponding to elevated p-MOS scores. Additionally, we utilize large language models (LLMs) and CLAP to synchronize music signals with captions, thereby enhancing text-audio correlation in extensive music datasets. Our ablation studies on public datasets confirm the effectiveness of our methodology, with the final model surpassing previous works in both objective and subjective measures.
2 Related Work
Text to music generation. Text-to-music generation aims to create music clips corresponding to input descriptive or summary text. Previous efforts have utilized either language models (LMs) or DMs to model quantized waveform representations or spectral features. For continuous waveform generation, MusicLM (Agostinelli et al., 2023), MusicGen (Copet et al., 2024), and MeLoDy (Lam et al., 2024) leverage LMs and DMs on residual codebooks obtained via quantization-based audio codecs (Zeghidour et al., 2021; Défossez et al., 2022). Moûsai (Schneider et al., 2023), Noise2Music (Huang et al., 2023), Riffusion (Forsgren and Martiros, 2022), and AudioLDM 2 (Liu et al., 2023a) use U-Net-related diffusion to model mel-spectrograms or latent representations obtained through a pretrained VAE, which are then converted into audio waveform via pretrained vocoders  (Kong et al., 2020). Existing methods lack consideration of low-quality data issues during the modeling phase and fail to effectively learn the distribution of high-quality audio. Our approach addresses this issue by introducing quality-aware training.
Transformer based diffusion models. Traditional diffusion models typically use U-Net as the backbone, where the inductive biases of CNNs do not effectively model the spatial correlations of signals and are insensitive to scaling laws (Li et al., 2024). However, these issues have been effectively addressed in transformer-based diffusion models (Peebles and Xie, 2023). This advantage has been particularly evident in fields such as video generation (Brooks et al., ), image generation (Peebles and Xie, 2023; Chen et al., 2024a; Bao et al., 2022), and speech generation (Liu et al., 2023b). To expedite training and foster inter-domain learning of correlations, the masking strategy has been proven effective (Lei et al., ), which has yielded the best class-conditioned results on ImageNet (Gao et al., 2023), while a simpler architecture (Zheng et al., 2023) incorporating reconstruction losses and unmasked fine-tuning further enhances model training speed. However, these models have not yet been verified for text-controlled audio generation capabilities on large-scale audio datasets, and their adaptability to additional control information remains an open question.
3 Preliminary
Latent diffusion model. Direct application of DMs to cope with distributions of raw signals incurs significant computational overhead Ho et al. (2020); Song et al. (2020). Conversely, studies Liu et al. (2023c, a) apply them in a latent space with fewer dimensions. The latent representation $z_{0}$ is the ultimate prediction target for DMs, which involve two key processes: diffusion and reverse processes. In the diffusion process, Gaussian noise is incrementally added to the original representation at each time step, described by $z_{t+1}=\sqrt{1-\beta_{t}}z_{t}+\sqrt{\beta_{t}}\epsilon$, where $\epsilon$ is drawn from a standard normal distribution $\mathcal{N}(0,I)$, and $\beta_{t}$ is gradually adapted based on a preset schedule to progressively introduce noise into the state $z_{t}$. The cost function Ho et al. (2020); Liu et al. (2023c) is formalized as:
$\arg\min_{\theta}\left[\mathbb{E}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(z_{0},y),\epsilon}}\left[\left\|\epsilon-D_{\theta}\left(\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,t,y\right)\right\|^{2}\right]\right],$ (1)
where $D_{\theta}$, the denoising model, strives to estimate the Gaussian noise $\epsilon$, conditioned on the latent state $z_{t}$, the time step $t$, the conditional embedding $y$, and where $\alpha_{t}$ represents a predefined monotonically increasing function. In the reverse process, we obtain $z_{t-1}$ via the recursive equation: $z_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(z_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}\right)+\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}}\epsilon$, where $\epsilon_{\theta}$ represents the estimated Gaussian noise.
Classifier-free guidance and negative guidance. Classifier-free guidance (CFG), introduced by (Ho et al., 2020), increases the versatility of DMs by enabling both conditional and unconditional generation. Typically, a diffusion model generates content based on specific control signals $y$ within its denoising function $D_{\theta}(z_{t},t,y)$. CFG enhances this mechanism by incorporating an unconditional mode $D_{\theta}(z_{t},t,\emptyset)$, where $\emptyset$ symbolizes the absence of specific control signals. The CFG-enhanced denoising function is then expressed as $D_{\theta}^{\text{CFG}}(z_{t},t,y)=D_{\theta}(z_{t},t,y)+w(D_{\theta}(z_{t},t,y)-D_{\theta}(z_{t},t,\emptyset))$, where $w\geq 1$ denotes the guidance scale. During training, the model substitutes $y$ with $\emptyset$ at a constant probability $p_{\text{uncond}}$. In inference, $y$ might be replaced by a negative prompt like "low quality" to prevent the model from producing such attributes (Liu et al., 2023a).
4 Method
Method overview. We adopt a three-tiered strategy to enable the proposed model to predict high-quality music waveform across a complex corpus. Initially, we introduce a waveform-quality-aware technique in Section 4.1 that enables the model to discern waveform quality, extracted from from both coarse and fine-grain perspectives. Next, in Section 4.2, we present our mask diffusion model to enhance the model’s capacity for quality waveform modeling and the management of conditional text information. Lastly, we outline a novel approach to address issues with problematic text, such as weak, inaccurate, or missing captions.
4.1 Quality information injection
At the heart of our work lies the implementation of a pseudo-MOS scoring model (Ragano et al., 2023) to meticulously assign music quality to quality prefixes and quality tokens.
We define our training set as $\mathcal{D}_{o}=\{(M_{i},T^{o}_{i})\mid i=1,2,\dots,N_{D}\}$, where each $M_{i}$ represents a music signal and $T^{o}_{i}$ is the corresponding original textual signal, we achieve the p-MOS set $S=\{s_{1},s_{2},\ldots,s_{n}\}$. To optimize model learning from datasets with diverse audio quality and minimize the impact of low-quality audio, we initially assign p-MOS scores to each music track using a model fine-tuned with wav2vec 2.0 (Baevski et al., 2020) on a dataset of vinyl recordings for audio quality assessment. These scores facilitate dual-perspective quality control for enhanced granularity and precision.
First, We analyze this p-MOS set $S$ to identify a negative skew normal distribution with mean $\mu$ and variance $\sigma^{2}$. We define text prefixes based on $s$ as follows: prepend "low quality" if $s<\mu-2\sigma$, "medium quality" if $\mu-\sigma\leq m\leq\mu+\sigma$, and "high quality" if $s>\mu+2\sigma$. This information is prepended before processing through the text encoder with cross-attention, enabling the initial separation of quality-related information.
To achieve a more precise awareness and control of waveform quality, we synergize the role of text control with quality embedding. We define the quantization function $Q:[0,5]\to\{1,2,3,4,5\}$, which maps p-MOS scores to discrete quality levels based on the distance from the mean $\mu$ in terms of standard deviation $\sigma$:
$Q(s)=\left\lfloor\frac{s-(\mu-2\sigma)}{\sigma}\right\rfloor+r$ (2)
where $r=2$ for $s>\mu$, otherwise, $r=1$. Subsequently, $Q(s)$ is mapped to a $d$-dimensional quality vector embedding using the embedding function $E$, such that
$q_{\text{vq}}(s)=E(Q(s))\in\mathbb{R}^{d},$ (3)
This process provides finer granularity of control within the following model and facilitates the ability of interpolative quality control during inference, enabling precise adjustments in $\mathbb{R}^{d}$.
Figure 2: Pipeline of proposed quality-aware masked diffusion transformer for music generation.
4.2 Quality-aware mask diffusion transformer
In a general patchify phrase with patch size $p_{f}\times p_{l}$ and overlap size $o_{f}\times o_{l}$, patchified token sequence $X=\{x_{1},x_{2},\dots,x_{P}\}\subset\mathbb{R}^{p_{f}\times p_{l}}$ are obtained through spliting the music latent space $\mathcal{M}_{spec}\in\mathbb{R}^{F\times L}$, as described in  5.2. The total number of patches $P$ is given by:
$P=\left\lceil\frac{L-p_{l}}{p_{l}-o_{l}}+1\right\rceil\times\left\lceil\frac{F-p_{f}}{p_{f}-o_{f}}+1\right\rceil$ (4)
A 2D-Rope position embedding (Su et al., 2024) is added to each patch for better modeling of relative position relationship while a binary mask $\boldsymbol{m}\in\{0,1\}^{P}$ is applied during the training stage, with a variable mask ratio $\gamma$. This results in a subset of $\lfloor\gamma P\rfloor$ patches being masked that $\sum_{i=1}^{P_{N}}m_{i}=\lfloor\gamma P\rfloor$, leaving $P-\lfloor\gamma P\rfloor$ patches unmasked. The subset of masked tokens is invisible in the encoder stage and replaced with trainable mask tokens in the decoder stage following the same strategy utilized in AudioMAE (Huang et al., 2022) and MDT (Gao et al., 2023).
The transformer we use consists of $N$ encoder blocks, $M$ decoder blocks, and an intermediate layer to replace the masked part with trainable parameters. We treat the embedding of the quantized p-MOS score as a prefix token, concatenated with each stage’s music tokens. Let ${X}^{k}=[{x}^{k}_{1},{x}^{k}_{2},\ldots,{x}^{k}_{P}]\in\mathbb{R}^{P\times d}$ represent the output of $k$-th encoder or decoder block, where the initial input of the encoder $X^{0}=z_{t}=\alpha_{t}{z}_{0}+\sqrt{1-\alpha_{t}}\epsilon$, and the final decoder block estimate $X^{N+M}=z_{0}={X}^{0}=[{x}_{1},{x}_{2},\ldots,{x}_{P}]$. For $k<N$, indicating the encoder blocks, the sequence transformation focuses only on unmasked tokens:
$[q_{\text{vq}}^{k+1};{X}^{k+1}]=\text{ Encoder}^{k}\left(\left[q_{\text{vq}};{X}^{k}\odot(\mathbf{1}-\boldsymbol{m})\right]\right),$ (5)
where $\boldsymbol{m}\in\{0,1\}^{P}$ is the mask vector, with $1$ indicating masked positions and 00 for visible tokens.
For $N<k<N+M$, indicating the decoder blocks, the full sequence including both unmasked tokens and learnable masked tokens is considered:
$[q_{\text{vq}}^{k+1};{X}^{k+1}]=\text{Decoder}^{k}\left(\left[q_{\text{vq}};{X}^{k}\right]\right),$ (6)
where the previously masked tokens are now subject to prediction and refinement. In the decoding phase, the portions that were masked are gradually predicted, and throughout this entire phase, the quality token $q_{\text{vq}}(s)$ is progressively infused and optimized. Subsequently, the split patches are unpatchified while the overlapped area is averaged to reconstruct the output noise and every token contributes to calculating the final loss:
$\mathcal{L}(\theta)=\mathbb{E}_{(z_{0},q_{vq},y),\epsilon}\left[\left\|\epsilon-D_{\theta}\left(\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,t,q_{vq},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y}\right)\right\|^{2}\right]$ (7)
In the inference stage, the model can be guided to generate high-quality music through CFG:
$D_{\theta}^{\text{High Quality}}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{high}}},y)=D_{\theta}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{high}}},y)+w(D_{\theta}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{high}}},y)-D_{\theta}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{low}}},\emptyset))$ (8)
Here $q_{vq}^{\text{high}}$ and $q_{vq}^{\text{low}}$ indicate quantified p-MOS for guiding the model in a balance between generation quality and diversity.
4.3 Music caption refinement
We divided the caption refinement stage into three steps including text information enriching with music caption model $\mathcal{F}_{\text{cap}}$, caption alignment adjustment with CLAP cosine similarity function $\mathcal{S}$ and caption diversity extension with LLMs which we denoted as ${\mathcal{F}}_{\text{llm}}$.
Initially, pretrained music caption model  (Doh et al., 2023) is employed to re-annotate each music signal, shown as $\mathcal{D}_{g}=\{(M_{i},T^{g}_{i})\mid T^{g}_{i}=\mathcal{F}_{\text{cap}}(M_{i}),i=1,2,\dots,N\}$. CLAP text-audio similarity is applied to filter $\mathcal{D}_{g}$ with a threshold of $\rho_{1}$, resulting in
$\mathcal{D}^{\text{filter}}_{g}=\{(M_{i},T^{g}_{i})\mid\mathcal{S}(T^{g}_{i},M_{i})>\rho_{1}\}$ (9)
In this context, we meticulously filter out generated captions that do not correspond with their respective audio files. This misalignment may be attributed to inaccuracies within the captioner’s insufficient training. For the filtered data, we opt to retain the use of the original captions.
To ensure that valuable information from the original captions is not overlooked when using only the generated captions, we first need to filter out original captions suitable for the fusing stage. Formulated as
$\mathcal{D}^{\text{filter}}_{o}=\{(M_{i},T^{o}_{i})\mid\mathcal{S}(T^{o}_{i},M_{i})>\rho_{2}\}.$ (10)
The issue can stem from the original data being improperly labeled with terms such as ’speech, car’ from datasets like AudioSet (Gemmeke et al., 2017) and also may be because of desperately missing of labels.
Finally, only the original caption that suffers low CLAP text similarity score should be merged with the generated ones, for redundant, repetitive parts result in long and verbose final captions. Thus, we set the threshold to $\rho_{3}$ and merge them by LLMs to $T_{\text{fusion}}=\mathcal{F}_{\text{llm}}(T^{o},T^{g})$:
$\mathcal{D}_{\text{merge}}=\left\{(M_{i},T_{\text{fusion}})\mid\mathcal{S}(T^{o},T^{g})<\rho_{3},(M_{i},T^{o})\in\mathcal{D}^{\text{filter}}_{o},(M_{i},T^{g})\in\mathcal{D}^{\text{filter}}_{g}\right\}.$ (11)
5 Experimental setup
5.1 Datasets
Train datasets. We used the following databases for our training: AudioSet Music Subset (ASM) (Gemmeke et al., 2017), MagnaTagTune (MTT) (Law et al., 2009), Million Song Dataset (MSD) (Bertin-Mahieux et al., 2011), Free Music Archive (FMA) (Defferrard et al., 2016), and a private dataset. Each track in these databases was clipped to 10-second segments and sampled at 16kHz to ensure uniformity across the dataset. The final training set was developed through a process of caption refinement, as detailed in Section 4.3. Finally, we got our training set totaling 12.5k hours of diverse music data. The specific composition of these datasets is further elaborated in the Appendix A.1.
Evaluation datasets. For comparison with prior work, we evaluate our model on the widely used MusicCaps benchmark (Agostinelli et al., 2023). MusicCaps consists of 5.5K $10.24$-second clips that include high-quality music descriptions from ten musicians.
5.2 Models and hyperparameters
Audio compression.
Each 10.24-second audio clip, sampled at 16 kHz, is initially transformed into a $64\times 1024$ mel-spectrogram with mel-bins of 64, hop-length of 160 and window length of 1024. Subsequently, this spectrogram is compressed into a $16\times 128$ latent representation $\mathcal{M}_{spec}$ using a Variational Autoencoder (VAE) pretrained with AudioLDM 2 (Liu et al., 2023a). This latent representation is utilized for adding noise and training purposes. Finally, we employ a pretrained Hifi-GAN (Kong et al., 2020) to reconstruct the waveform from the generated mel-spectrogram.
Caption processing and conditioning. We utilize the LP-MusicCaps (Doh et al., 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the LAION-CLAP (Wu* et al., 2023) checkpoint music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”. Do not add any imaginary elements and try to keep the modified caption refined and accurate. We use FLAN-T5-large (Peebles and Xie, 2023) as text encoder for all models.
Diffusion backbone. We train our diffusion model with three backbones for comparison: U-Net (Ronneberger et al., 2015) based at 1.0B parameters, code based on AudioLDM (Liu et al., 2023c); and our proposed Quality-aware Masked Diffusion Transformer based at 675M parameters with $N=20$ encoder layers and $M=8$ decoder layers. Flash attention (Dao et al., 2022) from the xFormers package (Lefaudeux et al., 2022) is utilized to improve training speed and memory usage for self-attention and cross-attention stages. We study the impact of the patch size and overlap size in the Appendix A.2 , and apply a patch size of $1\times 4$ without overlap for the training of our final model. We train on 10-second audio crops sampled at random from the full track, maintaining a total batch size of 64 on four NVIDIA A100 80GB GPUs, with gradient clipping of 1.0, learning rate of 8e-5, 2000 steps of linear warmup without weight decay, and a condition drop of 0.1 during training. The final model was trained for a total of seven days. During inference, we use Denoising Diffusion Implicit Models (DDIM)  (Song et al., 2020) with 200 steps and a guidance scale of 3.5, consistent with AudioLDM  (Liu et al., 2023c).
5.3 Evaluation metrics
We evaluate the proposed method using objective metrics, including the Fréchet Audio Distance (FAD) (Kilgour et al., 2018), Kullback-Leibler Divergence (KL), Inception Score (IS). All above metrics are computed using the audioldm_eval library (Liu et al., 2023c), ensuring standardized and reliable evaluation. We also utilize pseudo-MOS scoring model (Ragano et al., 2023) to estimate generation quality, with more accurate assessments derived from subjective metrics.
For the human studies, we ask human raters to evaluate two aspects of the audio samples (i) overall quality (Ovl), and (ii) relevance to the text input (Rel). For the overall quality test, raters were asked to rate the perceptual quality of the provided samples with 1 to 5. For the text relevance test, raters were asked to rate the match between audio and text on a scale from 1 to 5. Our evaluators consist of people from various backgrounds, including professional music producers, video editors who need to use soundtracks, staff responsible for the operations of music products and complete beginners who are unfamiliar with related knowledge. We evaluate randomly sampled file, where each sample was evaluated by at least 5 raters.
6 Results
We begin by presenting our approach to refining captioning, which includes the capability for quality awareness, transitioning from text-level control to token-level control. Finally, we compare proposed model with previous works on both subjective and objective metrics.
6.1 Impact of music caption refinement
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al., 2009). Table  1 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
Table 1: Comparison of model performance training on different textual representations, evaluated by FAD, KL(Sigmoid based), IS and CLAP score.
U-Net based MDT based
[…]
$\mathcal{D}_{o}$ 7.228 4.449 1.742 0.199 7.072 3.964 2.116 0.291
$\mathcal{D}_{g}$ 5.944 4.134 2.276 0.278 5.757 3.837 2.508 0.342
$\mathcal{D}_{\text{merge}}$ 5.867 4.132 2.294 0.284 5.637 3.804 2.625 0.350
From Table  1 we can also observe consistent trends: employing a captioner to transform audio annotations from sparse words into detailed sentences significantly improved the models’ generalization and diversity. This indicates that detailed annotations are essential for learning the relationship between the models and spectral features. Moreover, the filter and fusion stages led to enhancements across all metrics, highlighting the significance of precise, comprehensive annotations for generalization ability and control ability. We also found that compared to U-Net, the MDT architecture shows stable improvements in basic modeling metrics, making it a better backbone for music spectral modeling.
6.2 Quality awareness
Figure 3: Comparison of model performance under different quality prefixes on MTT-FS, while the blue dashed line represents the threshold set during training to distinguish the three quality prefix levels, and the red one represents the test set average p-MOS value.
This subsection aims to explore the effects and interactions of model control over quality tokens and quality text prefixes during the training phase, as well as their comparative effects across different models. In our previous MTT dataset of 1,000 test pairs, we filtered out pairs labeled with low quality or quality is poor to avoid confusion when acting with quality prefixes, resulting in a new subset of 519 entries, which we refer to as the MTT Filter Set (MTT-FS). Figure 3 shows the impact of different quality prefixes during inference when using quality as text prefix during training for U-Net and MDT based backbone. It is observed that U-Net, when infer with different quality prefixes, only shows minor changes in p-MOS scores and does not follow the threshold set for training. In contrast, MDT is better at learning quality information from prefixes, achieving p-MOS scores far above U-Net and the test set. Additionally, through decoupling of quality information from training set, we earn superior diversity and generalizability than training and inference without quality text prefix.
Figure 4: (Left) Five VQ-MOS distribution curves are obtained by concurrently using text quality prefixes and quality tokens as controls on the MTT-FS, with quantized MOS levels ranging from 1 to 5 serving as control constraint inferences. The distribution of the training set is normalized by each sample’s duration, colored lines represent thresholds of quantized p-MOS tokens during training. (Right) The effect of using quality text prefixes during training is shown, showcasing testing results on four metrics: FAD, KL, IS, and p-MOS, while gray lines for quantized p-MOS.
Given that quality tokens are specifically designed for the Transformer architecture, Figure 4 (left) demonstrates the controlled outcomes when different quality tokens are employed after integrating quantified quality as a token during training. Remarkably, using quality tokens alone provided more precise and accurate p-MOS score control.
For our ablation study, we compared the effects of solely using text prefixes against combining both approaches. As shown in Figure 4 (right), with given quantized control level increasing gradually, the model steadily improves p-MOS scores, which represent the quality of generation. Concurrently, FAD and KL also optimize progressively until a turning point occurs at level 4 while higher average p-MOS is reached than sorely using prefix. The turning point may be due to the scarcity of examples with quality level 5 in the dataset. Additionally, through combining two types of quality information injection, the refined decoupling and interaction allowed the model to more accurately perceive audio data quality features through training, leading to significantly reductions in FAD and KL compared to only using either of them. We also compare our approach with the traditional "Negative Prompt" strategy in Appendix A.3, highlighting our approach’s significant improvement in quality and reduction in FAD.
6.3 Compared with previous methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al., 2023a) and MusicLDM (Chen et al., 2024b), which model music using spectral latent spaces, and MusicLM (Agostinelli et al., 2023), MeLoDy (Lam et al., 2024), and MusicGen Copet et al. (2024), which focus on modeling discrete representations. Additionally, we train AudioLDM (Liu et al., 2023c) on our own dataset with quality prefix for comparing between backbone designing and the use of quality token.
Table 2: Objective evaluation results for music generation with diffusion based models and language model based models. We re-infer MusicLDM and AudioLDM 2 with label †.
Details Objective
Model Params Hours Fad $\downarrow$ Kl $\downarrow$ Is $\uparrow$ Clap $\uparrow$
MusicLM 1290M 280k 4.00 - - -
MusicGen 1.5B 20k 3.80 1.22 - -
Mousai 1042M 2.5k 7.50 1.59 - -
MeLoDy 726M 257.0k 5.41 - - -
MusicLDM† - 455.6 5.35 2.54 2.26 -
AudioLDM 2 - Full 712M 17.9k 3.13 1.20 - -
AudioLDM 2 - Music † 712M 10.8k 4.04 1.46 2.67 0.34
Ours (U-Net) 1.0B 12.5k 2.03 1.51 2.41 0.33
Ours (QA-MDT) 675M 12.5k 1.65 1.31 2.80 0.35
For AudioLDM 2, MusicLDM, we re-inferred based on their official links and checkpoints to compare additional metrics under the same environment. The results are shown in Table  2. Based on the experimental results, we achieved significant advantages in both subjective and objective metrics. Since the Kullback-Leibler divergence is used to measure the distance between audio samples, higher quality audio usually results in deviations from the original waveform of the dataset, thereby leading to lower performance on the KL metric.
Table 3: Evaluation of model performances among different groups, rated for text relevance (Rel) and overall quality (Ovl), with higher scores indicating better performance. The groups included Production Operators (Po), Professional Music Producers (Pmp), and Video Editors (Ve).
Po Pmp Ve Beginners Average
Model Ovl Rel Ovl Rel Ovl Rel Ovl Rel Ovl Rel
Ground Truth 4.00 4.00 4.47 3.60 4.10 3.80 3.87 3.87 3.72 3.60
AudioLDM 2 2.03 2.42 3.03 3.61 3.21 3.71 3.85 3.85 3.03 3.40
Ours(U-Net) 2.80 3.34 3.46 4.08 3.40 3.96 3.88 3.96 3.38 3.84
Ours(QA-MDT) 3.27 3.77 3.69 4.19 3.54 3.94 4.23 4.00 3.68 3.97
Based on subjective evaluation shown in Table 3, our proposed method significantly improves overall audio quality and text alignment, thanks to the label optimization scheme for large music datasets and the quality-aware training strategy. By analyzing the backgrounds of the evaluators and their corresponding results, we can also see that for beginners, the comparison between different systems is not sensitive, which is related to their lack of music background experience and knowledge. However, from the perspective of our method in product operators, video editors, and audio producers, our method offers considerable enhancements, underscoring its potential value to audio industry professionals.
7 Conclusion and discussion
In this study, we identify the challenges posed by large-scale uneven audio quality and unaligned textual annotations in the music generation domain, which hinder the development of diffusion-based TTM generation. By employing a novel quality decoupling learning approach based on p-MOS, along with masked diffusion transformer as the backbone for the diffusion process, we achieve enhanced generation quality and musicality for music generation.
Limitations.
Music, as a form of art, requires consideration of aesthetic qualities in generated melodies, in addition to quality. In this work, we did not explicitly optimize the generated melodies, nor did we explore the extrapolative performance of transformers on long-duration audio generation, which we will leave as future work.
Broader impact.
Except for the scaling ability, transformer-based generative models offer greater flexibility compared to U-Net, and the mask strategy facilitates tasks such as music inpainting and super-resolution spectral expansion. We hope our work paves the way for advancements in music generation, control, and other related applications.
Appendix A Appendix
A.1 Training dataset.
Table A.1: Database statistics.
Database Duration Ave Length Clip Num Sample Rate Released
MTT 200h 30s 24K - 2009
FMA 900h 30s 10.9K 44.1kHz 2016
MSD 7333h 30s 880K 32kHz 2011
ASM 2777h 10s 1.0M mix 2017
Private 1375h 90s 55K mix -
Final 12.5kh 10s 45M 16kHz -
A.2 Ablation study for MDT structure
In image generation, for latent spaces with consistent aspect ratios, the patchify approach indicates that smaller square patch sizes ($2\times 2$ and $1\times 1$) achieve the best results [Peebles and Xie, 2023]. However, for our audio latent space, which has a size of $16\times 256$ with significant aspect ratio imbalance, the traditional square patch pattern may not be the best choice. Therefore, we chose $2\times 4$ as the fundamental patch size and conducted comparisons from three aspects: the decoder layer, patchify strategy, and in contrast to the traditional DiT.
Due to the potential inaccuracies inherent in tags, we evaluate our model on the Generated Captions without Filter and Fusion to assess its robustness against possibly inaccurate labels training with ASM and FMA, and test on MTT Test Set.
Decoder Layers
Table A.2: Performance comparison between numbers of decoder layers in the MTT Test Set.
Layers FAD $\downarrow$ KL $\downarrow$
6 5.937 3.912
8 5.901 3.912
10 5.958 3.724
To avoid introducing changes in the model’s parameter count, we fixed the total number of layers in the Encoder and Decoder at 28. As shown in the Table A.2, as the number of Decoder layers increases, the generality peaks at 8, and the sound quality gradually improves while diversity decreases. By conducting listening tests on a small subset of randomly sampled datasets, we found that the decrease in diversity leads to increasingly chaotic generated melodies, causing discomfort in the listening experience. Moreover, considering we have additional methods to control the quality of sound generation, to balance various metrics, we chose a layer count of 8 for subsequent experiments.
Pachify Strategy
In this section, we aim to explore the fundamental impact of different basic modeling units on spectral. From the experimental comparison of patch sizes $2\times 4$, $1\times 4$, and $2\times 2$ shown in Table A.3, it is evident that reducing the patch size consistently leads to performance improvements due to the more detailed spectral modeling. Considering the inherent spatial correlations in the Fourier transform within music spectrum, we analyze the results of experiments that apply spectral overlaps of 2 and 1 in the time and frequency domains separately, chich indicates that introducing overlap in the latent space does indeed contribute to improved results.
Table A.3: Performance comparison between different settings of patchify strategies on MTT Test Set.
Model Patch Size Overlap Size FAD $\downarrow$ KL $\downarrow$
DiT $2\times 4$ $0\times 0$ 6.861 4.355
MDT $2\times 4$ $0\times 0$ 5.901 3.913
MDT $2\times 2$ $0\times 0$ 5.685 3.820
MDT $1\times 4$ $0\times 0$ 5.768 3.837
MDT $2\times 4$ $1\times 0$ 5.757 3.837
MDT $2\times 4$ $0\times 2$ 5.172 3.737
A smaller patch sizes, such as $2\times 1$ or even the more extreme $1\times 1$, might offer improved performance. However, due to the significantly increased training and inference costs that accompany these sizes, we refrained from conducting further detailed experiments. Additionally, we explored the impact of these settings in comparison with DiT [Peebles and Xie, 2023, Chen et al., 2024a] architectures and discovered that the mask strategy, leveraging the spectral correlations in both the frequency and time domains, markedly enhances spectral modeling, which is evident both in terms of the convergence rate and the quality of the final results. However, based on subjective listening on a small test dataset, we find that although the overlap strategy can significantly improve the model’s objective indicators, it leads to a decline in both melodic generation and aesthetic aspects, which is crucial for music generation. Therefore, we choose $4\times 1$ as the base pachify strategy for subsequent experiments and the final model.
A.3 Comparing with using "negative prompt"
In this section, we aim to compare and analyze the traditional "negative prompt" method with our approach. In the implementation of CFG, we use
$D_{\theta}^{\text{CFG}}(x,t,y)=D_{\theta}(x,t,y)+w\left(D_{\theta}(x,t,y)-D_{\theta}(x,t,y_{neg})\right),$ (12)
where $y_{neg}$ is formulated as text embedding of "low quality."
Table A.4: Performance comparison between three systems on MTT-FS.
System FAD $\downarrow$ KL $\downarrow$ p-MOS $\uparrow$
MDT 5.757 3.837 3.796
MDT + Negative prompt 5.641 3.461 3.832
QA-MDT 5.200 3.214 4.051
Based on the Table A.4, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of "low quality" in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of "low quality" might not be well disentangled during training, leading to an suboptimal results.